Site 1166 | Table of Contents

SITE 1165

Hole 1165A (Proposed Site PBD-12A)
A beacon was dropped on the Global Positioning System (GPS) coordinates of Site 1165 at 1500 hr 28 January 2000. Once the hydrophones and thrusters were deployed and the vessel settled on location, the corrected precision depth recorder depth referenced to the dual elevator stool indicated a seafloor depth of 3552.4 meters below rig floor, equivalent to 3541.3 meters below sea level (mbsl).

Hole 1165A was spudded using the advanced hydraulic piston corer (APC) at 0450 hr on 29 January 2000. The hole was scheduled for a single mudline core, which recovered 5.43 m of sediment and placed the drill pipe-measured seafloor depth at 3537.0 mbsl. The bit was pulled clear of the mudline, and the hole was abandoned at 0515 hr.

Hole 1165B
An offset was not deemed necessary, and Hole 1165B was spudded at 0610 hr on 29 January 2000. The APC bit was positioned at 3534.95 mbsl and the water depth was established at 3537.65 mbsl, based on recovery of the mudline core. APC coring continued through 18H to a depth of 148.3 mbsf. Piston coring was abandoned at that depth because of a continuing problem with jammed liners and poor recovery. Adara temperature tool measurements were taken at the mudline and at the depth of Cores 4H, 7H, 10H, and 14H. Data from the position of Core 7H could not be retrieved. Coring proceeded using the extended core barrel (XCB). Operations were suspended when iceberg Nora approached to within 2.0 nmi of the drill site. Core 29X was recovered from a depth of 252.4 mbsf, and the drill string was pulled up to 50 mbsf. At 1630 hr on 30 January 2000, the iceberg had moved to a distance of 2.6 nmi on a course that was taking it directly away from the drill ship. XCB coring resumed at 1800 hr on 30 January. Coring proceeded with excellent results through Core 67X to a depth of 607.3 mbsf. The XCB system was pushed in hard material beyond its normal use to a depth of 682.2 mbsf (Core 76X) because it was assumed that the hard formation might be penetrated in a few tens of meters. Unfortunately, the steadily declining rate of penetration and poor core recovery indicated that this was not the case. The last XCB core was recovered on 3 February at 1500 hr, and the decision was made to terminate coring in Hole 1165B in favor of rotary core barrel (RCB) coring. The hole was displaced with 18.8 bbl of cement, and the pipe was tripped back to the surface. The rotary table was cleared at 0530 hr on 4 February, ending Hole 1165B.

Hole 1165C
The vessel was offset 50 m to the west, and Hole 1165C was spudded at 1805 hr on 4 February. The seafloor depth was established as 3537.55 mbsl. The hole was washed down to a depth of 54.0 mbsf, where Core 1R was taken from 54.0 to 63.6 mbsf, recovering an interval missed in Hole 1165B. A center bit was deployed, and the hole was drilled down to a depth of 673.0 mbsf. Continuous RCB coring began at that depth and continued until 1445 hr on 6 February, when iceberg Mona came to within 4 nmi of the drill site. Core 5R was recovered, the drill pipe was pulled back to a depth of 75.1 mbsf, and a free-fall funnel (FFF) was deployed.

Mona approached within 2.1 nmi before moving on a path away from the ship. The pipe was tripped back in the hole at 0500 hr on 7 February, and RCB coring resumed at 1015 hr. A total of 19.25 hr was lost as a result of the approach of iceberg Mona. Continuous RCB coring continued until iceberg Lea headed toward the drill site on a southeasterly course. At 2015 hr on 8 February, coring had to be suspended after recovering Core 13R from a depth of 893.6 mbsf. Lea was moving rapidly toward the drill site at a speed of 0.7 kt and with a closest point of approach of 1.0 to 3.0 nmi, depending on the assumed course. The drill string was pulled to a depth of 75.1 mbsf. At 0600 hr on 8 February, the iceberg reversed course to north-northwest and sped up to 0.3 kt. Lea appeared to be moving away by 1115 hr on 8 February and was no longer considered a threat to drilling operations. The drilling assembly was run to bottom, and at 1515 hr on 9 February continuous RCB coring resumed. A total of 19.0 hr was lost due to the first encounter with iceberg Lea.

Coring proceeded with excellent results until Lea turned around and headed back toward the location, closing to within 5.0 nmi. Coring was suspended with the recovery of Core 32R from a depth of 970.2 mbsf, and the drill string was pulled to a depth of 75.1 mbsf. Lea moved to within 0.5 nmi of the ship's position at 2015 hr on a course and bearing likely to bring it even closer to the drillship. The decision was made to pull the drill pipe clear of the seafloor and to move the ship 0.7 nmi away from the location. Lea passed within 200 m of Site 1165. Ultimately Lea moved away, and the ship moved back over the location. Hole 1165C was reentered at 0250 hr on 11 February, the pipe was run to bottom, and coring resumed by 0915 hr. A total of 25.0 hr was lost as a result of the second encounter with iceberg Lea.

Only one core (33R) was cut before another iceberg, Bertha, approached to within 5.0 nmi of Site 1165. The drill string was pulled to the surface for the third time. Before Bertha moved out of range, another smaller, apparently wind-driven iceberg, Bertie, moved into the safety zone. At 1145 hr on 12 February, Bertie was at 3.7 nmi away from the location and moving away. The drill string was run back into the hole. A total of 28.0 hr was lost because of icebergs Bertha and Bertie.

Cores 34R and 35R were recovered before coring ended at a total depth of 999.1 mbsf. The pipe was raised to 89.7 mbsf, and preparations for logging were made.

Logging Operations in Hole 1165C
The first suite of logging tools was ready to be deployed at 0915 hr on 13 February. A 2-hr interruption was caused by an iceberg that was headed to the location from a distance of 3.9 nmi but changed course abruptly away from the drill site. The first logging string, the triple combination tool (triple combo), was deployed at 1200 hr. The triple combo string was composed of the dual-induction tool model E (DITE), high-temperature lithodensity sonde (HLDS), neutron array porosity sonde (APS), and high-temperature natural gamma sonde (HNGS). Unfortunately, the winch operator was unable to lower the tools past 118.3 mbsf. All attempts to pass beyond the tight spot or ledge were futile, and the decision was made to recover the logging tools and lower the drill string three additional stands to 175.9 mbsf. This time the tools reached 991.3 mbsf, or 7.8 m above the total depth of the hole. Good logs were recovered from this run, and the tools were recovered at 2400 hr.

The tools for the second logging run consisted of the Formation MicroScanner (FMS), digital sonic imager, and natural gamma tool. This string was run in the hole at 0215 hr on 14 February; however, a restriction in the hole prevented passage beyond 580.3 mbsf. A complete hardware failure prevented any FMS data from being collected. Sonic and natural gamma data were successfully obtained from the upper part of the hole. Hole 1165C was abandoned by setting a 30-m balanced plug cement.

The pipe trip from the seafloor commenced at 1315 hr, and by 2030 hr on 14 February the rig floor was secured, the thrusters were raised, the beacons were recovered, and the ship was under way for Site 1166.

Transit to Site 1166 (PBS-9B)
Based on satellite images from the National Ice Center (Washington, D.C.) and reports we had received from the Hakurei-Maru, a research vessel that had been operating in the area a week before our approach, it was unlikely that the primary shelf site (PBS-2A) would be ice free. Our plan was to locate the shelf site as far to the west as the ice would allow by heading for proposed site PBS-9A and proceed from there to the west toward proposed sites PBS-8A, PBS-7A, and possibly PBS-1A. During the transit into Prydz Bay, with ice floes to either side of the vessel, it became clear that it was unlikely that sites farther to the west would be ice free enough for drilling. The decision was therefore made to drill at site PBS-9A. Unfortunately, we discovered that a large iceberg was located directly over the selected drill site. As we were rapidly approaching our survey way point, we decided to move the site 3 nmi to the northwest. Permission for the new drilling location, dubbed PBS-9B, was requested by telephone and fax and granted during the approach to the site. The site survey ended before reaching the drill site because one water gun froze up and the other water gun broke an air hose. There was not enough time to fix these problems because the ship had to be on location before dark, in view of the close proximity of numerous icebergs and ice floes in the area.

Site 1166 | Table of Contents